634b87d1572e0aabf0c077b8a3c0e098.ppt
- Количество слайдов: 14
Project report
Status: MICE Muon Beam • Target: – Target presently installed ‘in-operative’: • Optical connections removed (why? ) – Implies realignment/set-up – Jacking motor drive/control not implemented • Implies remove target as soon as possible: – To recover some components if possible – To prepare for installation of final target • Plan being prepared – Copper and optical cables procured: • Need to be installed when available • Plan being prepared – Signiticant progress on bearings: • Brass and copper bearings show less wear • Diamond-like carbon on ceramic to be tested • Working towards documented solution to be presented to ISIS – Targets: • Prepare new target installation in November short shutdown
Status: MICE Muon Beam • Upstream beam line (Krendler et al): – Quad triplet and dipole 1 installed • Services (power and water) yet to be connected – Issues: • Vacuum box for D 1: – Possible need to reposition (to be studied) • Vacuum box for Q 1— 3: – Too big! (2 mm in diameter) – New box will be procured (Krendler) and installed – Magnetic testing (Tilley) • Dipole and all quads tested (implies water connection and power) and fields measured prior to installation – All magnets checked out OK
Status: MICE Muon Beam • PSI solenoid (Courthold et al. ): – Vacuum leak traced: • ‘Beam degrader’ was located in bore. Had been cooled using LHe. When degrader was removed, pipes had been cut and not re-sealed, no information passed on • W. Gloor (PSI) provided rails to draw out cold mass, visitted RAL to work with local team • Circuits tested under pressure and under vacuum – Major milestone! • But, many tests still to do
Status: MICE Muon Beam • Linde refrigerator (Courthold et al): – Intend to start installation commissioning 20 Aug 07 – Many hurdles overcome: • Bureaucratic – CDM/certification issues now final step: • Propose to meet with Jonathan Carkeet
Status: MICE Muon Beam • Plinth (Jones et al): – Steel boxes installed and aligned – Concrete fill complete – Plates installed on top, levelled – Grouting complete – First layer of rail system installed and aligned – Next steps on installation awaits completion of shielding walls
Status: MICE Muon Beam • Survey: – Synchrotron-side survey complete (Loughrie) – MICE Hall side survey partially complete (Jones, Croxford): • Beam-line out of synchrtron, through PSI solenoid, D 2, and into Hall surveyed • Preliminary survey of D 2 rails performed: – Need more detailed survey to define shimming • Survey of plinth rail system needed as it is installed – Required equipment on ‘the right side of the shielding’!
Status: MICE Muon Beam • Dipole 2 (Jones et al): – Rail system installed – Bogey installed on rails – D 2 installed on bogey in ‘out’ position – Issues: • Require to design/implement movement system • Services (power/water) to be installed • Magnetic measurements to be made
Status: MICE Muon Beam • Q 35 refurbishment (for Q 4—Q 9) (Eammon, Cyril): – First Q 35 reassembled • Final steps in re-fitting of water manifold being carried out now – Schedule: • To be revised now that first complete magnet is available
Status: MICE Muon Beam • Shielding: – Removable north wall: • Complete – South wall steel and bracing • Complete – Synchtron (east) wall concrete infill • Complete – South wall concrete block/fill wall • In progress – Down stream (west) wall steel • In progress – Down stream (west) concrete • Reinforced concrete blocks for final build on order – Annoying, but no real impact on programme • Temporary concrete blocks available and will be installed this week (perhaps at the weekend)
Status: MICE Hall • House keeping: – – CMS equipment (magnet etc) removed CMS cabin demolished Q 4 s removed Tidying: • Much tidying done, still quite a lot more to do! • Infrastructure and services: – Water (Governs): • Preparing to procure system and install to be available 01 Dec 07 – Air: • Reconsider heat load into air: – Options to air-condition available – Magnetic shielding • Final step: specification of permittivity – Floor plates etc. • Designs available, need to push through – Urgent to move on the design/implementation as soon as shielding build complete
Status: MICE instrumentation • Upstream PId: – – TOF 0/1: on track for delivery to RAL ~Nov CKOV: Parts of first module now in R 12 • Mirror production at FNAL • • Second module to be shipped Preparing to take part in cosmic test with tracker starting September/October – • Spectrometer solenoid: – – • Slight delay Fabrication: steady progress Expect first solenoid at RAL in February Tracker: – Further 2 VLPC cassettes released to MICE by D 0 – Electronics: • Important milestone, now have 75% of what we need • VLSB: some uncertainty owing to internal re-organisation at FNAL • Firmware development: satisfactory progress – – Mechanical: • • Some delay, casualty of pressure on beam line (Barber) Station acceptance being brought up at Imperial All stations for both trackers available To do: – – • • Probably OK, already enough boards for cosmic test Detailing of final supports Completion of model, including diffuser Cabling to be prepared at FNAL Cosmic rig design Able to move this up the priority list now Downstream PId: – TOF 2, CAL • Need to understand how these will be provided
Status: RF, LH 2, FCs, RFCC • RF: – Power out of 250 k. W amplifier in June – Plan for power out of MW amplify by Christmas • • Need to check LH 2: – Cryostat being procured – Issues to be address: • • Minor modifications to design to be reviewed Safety to be reviewed – • Papers in preparations » TB will take appropriate action FC modules: – Tender process underway: • • Plan to close on schedule in September RFCC: – Sorry, assumed MZ would make comments • C&S: – First round of preparation done at June collaboration meeting – Need now to combine schedules/cost estimates – Need to make progress on this
Schedules: • Casualty of disc crash! • Apologies – Preliminary schedule for December shutdown prepared: • Critical – Schedule for work August—December in preparation • Again critical – Schedule for completion of Phase I and Phase II • Urgent • Will pick this up asap
634b87d1572e0aabf0c077b8a3c0e098.ppt